OpenStudy (logan13):

What are the solutions to the equation x² = 256? Choose all answers that are correct. A. 128 B. –128 C. 16 D. –16

OpenStudy (anonymous):

C and D

OpenStudy (anonymous):

16*16=256 -16*-16=256